ebhat is going on vacation in exactly 4 months. His options for paying for it are
1. pay $4000.00 up front (today)
2. make 4 payments of $X at the end of each month for the next 4 months.
3. Pay $1000.00 today and pay 3 * $X on the day he leaves (t = 4 months).
a) If X = 1100.00 what monthly compounded interest rate r(12) makes options 1 and 2 equally expensive?
b) Use your answer to a) to find the cost (present value) of option 3).
c) True or False: Option 3 is always cheaper than Option 2 no matter what the interest rate is.
1
Expert's answer
2020-06-15T17:45:48-0400




"1100\u00d74=4400"

"P(1+r)^n"

"4000(1+r)^4=4400"

"(1+r)^4=1.1"

"1+r=1.02411"

"r=0.02411=2.411\%"

The compound intrests that makes a more expensive should be >2.411%

b.The future cost of this amount is

"1000+(3\u00d71100)=4300"

Pv of a future sum =


"Fv\over(1+r)^n"

"4300\over(1+{2.411\over100})^4"

"=3909.0923"

c .Its true that option 3 is cheaper than option 2